# Theoretical Insights into Overparameterized Models in Multi-Task and Replay-Based Continual Learning

### v1

@@ C.1 Useful Lemmas
    Without loss of generality, we can focus on finding $\mathbb{E}[\|P_{1}u\|^{2}]$ where $u\in\mathbb{R}^{p}$ and $\|u\|^{2}=1$, since $X_{1}$ is a Gaussian matrix with rank $n_{1}$, and $P_{1}=X_{1}(X_{1}^{T}X_{1})^{-1}X^{T}_{1}$ is an orthogonal projection matrix with a similar rank. Due to the rotational invariance of the standard normal distribution, we can assume that $P_{1}$ is a fixed matrix and instead, $u$ is a random vector uniformly sampled from the unit sphere in $\mathbb{R}^{p}$. Using the rotational invariance again, we may assume without the loss of generality that $P_{1}$ is the coordinate projection onto the first $n_{1}$ coordinates in $\mathbb{R}^{p}$. Then it holds that:
@@
    $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}[\operatorname{tr}(w^{T}P_{1}X_{2}(X_{2}^{T}(I-P_{1})X_{2})^{-1}X^{T}_{2}P_{1}w)]$
    $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\operatorname{tr}(\mathbb{E}_{X_{2}}[X^{T}_{2}P_{1}ww^{T}P_{1}X_{2}(X_{2}^{T}(I-P_{1})X_{2})^{-1}])]$
@@
    $\Rightarrow\mathbb{E}[\|P_{01}w\|^{2}]$ $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\operatorname{tr}(\mathbb{E}_{X_{2}}[X^{T}_{2}P_{1}ww^{T}P_{1}X_{2}]\mathbb{E}_{X_{2}}[(X_{2}^{T}(I-P_{1})X_{2})^{-1}])]$
    $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\operatorname{tr}(\operatorname{tr}(P_{1}ww^{T}P_{1})\mathbb{E}_{X_{2}}[(X_{2}^{T}(I-P_{1})X_{2})^{-1}])]$
    $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\|P_{1}w\|^{2}\operatorname{tr}(\mathbb{E}_{X_{2}}[(X_{2}^{T}(I-P_{1})X_{2})^{-1}])]$
    We focus on finding the inner expectation first. Notice that for a fixed $X_{1}$, $I-P_{1}$ is an orthogonal projection matrix with rank $p-n_{1}$. Due to the rotational invariance of the standard normal distribution, we may assume without loss of generality that $I-P_{1}$ is the projection matrix that projects onto the first $p-n_{1}$ coordinates. With this in mind, $(X_{2}^{T}(I-P_{1})X_{2})^{-1}$ follows an inverse-Wishart distribution with an identity scale matrix $I_{n_{2}\times n_{2}}\in\mathbb{R}^{n_{2}}$ and $p-n_{1}$ degrees of freedom.
    $\Rightarrow\mathbb{E}[\|P_{01}w\|^{2}]$ $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\|P_{1}w\|^{2}\operatorname{tr}(\frac{I_{n_{2}\times n_{2}}}{p-n_{1}-n_{2}-1})]$
@@
Lemma C.8.
Assume matrices $X_{1}\in\mathbb{R}^{p\times n_{1}}$, $X_{2}\in\mathbb{R}^{p\times n_{2}}$ and $X_{3}\in\mathbb{R}^{p\times n_{3}}$ to be random matrices with entries being independently sampled from the standard Gaussian distribution $\mathcal{N}(0,1)$. Assume $p>n_{1}+n_{2}+n_{3}+1$ and let $X_{13}\in\mathbb{R}^{p\times(n_{1}+n_{2}+n_{3})}$ be the result of concatenation of $X_{1}$, $X_{2}$, $X_{3}$. Also let $X_{01},X_{02},X_{03}\in\mathbb{R}^{p\times(n_{1}+n_{2}+n_{3})}$ be the result of concatenating $X_{1}$, $X_{2}$ and $X_{3}$ with zero matrices such that $X_{01}=\begin{bmatrix}X_{1}&0&0\end{bmatrix}$ , $X_{02}=\begin{bmatrix}0&X_{2}&0\end{bmatrix}$ and $X_{03}=\begin{bmatrix}0&0&X_{3}\end{bmatrix}$. Define $P_{01}=X_{13}(X^{T}_{13}X_{13})^{-1}X_{01}^{T}$ and $P_{03}=X_{13}(X^{T}_{13}X_{13})^{-1}X_{03}^{T}$. Assuming $w,w^{\prime}\in\mathbb{R}^{p}$ are fixed given vectors, it holds that
$\mathbb{E}[w^{T}P^{T}_{01}P_{03}w^{\prime}]=-\frac{n_{1}n_{3}}{p(p-(n_{1}+n_{2}+n_{3})-1)}\langle w,w^{\prime}\rangle$
@@
Based on Lemma C.4, we have
$\mathbb{E}[w^{T}P^{T}_{01}P_{03}w^{\prime}]$ $=-\mathbb{E}[w^{T}P_{1}(I-\hat{P}_{12})X_{3}(X_{3}^{T}(I-P_{1})(I-\hat{P}_{12})X_{3})^{-1}X^{T}_{3}w^{\prime}]$
$=-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[w^{T}P_{1}(I-\hat{P}_{12})X_{3}(X_{3}^{T}(I-P_{1})(I-\hat{P}_{12})X_{3})^{-1}X^{T}_{3}w^{\prime}]]$
where $\hat{P}_{12}=X_{2}(X^{T}_{2}(I-P_{1})X_{2})^{-1}X^{T}_{2}(I-P_{1})$. Now denote $P=(I-P_{1})(I-\hat{P}_{12})$. With a fixed $X_{1}$ and $X_{2}$, $P$ is an orthogonal projection matrix, meaning that $P^{T}=P^{2}=P$. Therefore,
$\mathbb{E}[w^{T}P^{T}_{01}P_{03}w^{\prime}]$ $=-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[w^{T}P_{1}(I-\hat{P}_{12})X_{3}(X_{3}^{T}PX_{3})^{-1}X^{T}_{3}(P+I-P)w^{\prime}]]$
$=-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[w^{T}P_{1}(I-\hat{P}_{12})X_{3}(X_{3}^{T}PX_{3})^{-1}X^{T}_{3}Pw^{\prime}]]$
$-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[w^{T}P_{1}(I-\hat{P}_{12})X_{3}(X_{3}^{T}PX_{3})^{-1}X^{T}_{3}(I-P)w^{\prime}]]$
@@
$P_{1}(I-\hat{P}_{12})X_{3}$ $\operatorname{\perp\!\!\!\perp}(I-P_{1})(I-\hat{P}_{12})X_{3}$
@@
$\mathbb{E}[w^{T}P^{T}_{01}P_{03}w^{\prime}]$ $=-\mathbb{E}_{X_{1},X_{2}}[w^{T}P_{1}\mathbb{E}_{X_{3}}[(I-\hat{P}_{12})X_{3}]\mathbb{E}_{X_{3}}[(X_{3}^{T}PX_{3})^{-1}X^{T}_{3}Pw^{\prime}]]$
$-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[\operatorname{tr}(X^{T}_{3}(I-P)w^{\prime}w^{T}P_{1}(I-\hat{P}_{12})X_{3}(X_{3}^{T}PX_{3})^{-1}])]$
$=0-\mathbb{E}_{X_{1},X_{2}}[\operatorname{tr}(\mathbb{E}_{X_{3}}[X^{T}_{3}(I-P)w^{\prime}w^{T}P_{1}(I-\hat{P}_{12})X_{3}]\mathbb{E}_{X_{3}}[(X_{3}^{T}PX_{3})^{-1}])]$
$=-\mathbb{E}_{X_{1},X_{2}}[\operatorname{tr}((I-P)w^{\prime}w^{T}P_{1}(I-\hat{P}_{12}))\operatorname{tr}(\mathbb{E}_{X_{3}}[(X_{3}^{T}PX_{3})^{-1}])]$
Here we again use a technique similar to the one we used in Lemma C.7 part (iii) by first focusing on the inner expectation. Assume a fixed $X_{1}$ and $X_{2}$. Then $I-P_{1}$ and $P_{1}$ are orthogonal projections with ranks $p-n_{1}$ and $n_{1}$ respectively. Additionally, $(I-P_{1})\hat{P}_{12}$ is an orthogonal projection matrix with rank $n_{2}$ and $p-n_{1}>n_{2}$. On the other hand, $P_{1}$ and $(I-P_{1})\hat{P}_{12}$ are orthogonal projections. Therefore, $P_{1}+(I-P_{1})\hat{P}_{12}$ is a projection matrix with rank $n_{1}+n_{2}$ and $P=I-(P_{1}+(I-P_{1})\hat{P}_{12})$ is a projection matrix with rank $p-(n_{1}+n_{2})$. When taking the inner expectation only w.r.t. $X_{3}$, due to the rotational invariance property, we may assume that $(I-P_{1})(I-\hat{P}_{12})$ is the coordinate projection onto the first $p-(n_{1}+n_{2})$ coordinates. Therefore, $(X_{3}^{T}(I-P_{1})(I-\hat{P}_{12})X_{3})^{-1}\sim\mathcal{IW}(I_{n_{3}\times n_{3}},p-(n_{1}+n_{2}))$ and we have:
$\Rightarrow\mathbb{E}[w^{T}P^{T}_{01}P_{03}w^{\prime}]$ $=-\mathbb{E}_{X_{1},X_{2}}[\operatorname{tr}((I-P)w^{\prime}w^{T}P_{1}(I-\hat{P}_{12}))\operatorname{tr}(\frac{I_{n_{3}\times n_{3}}}{p-(n_{1}+n_{2})-n_{3}-1})]$
$=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}_{X_{1},X_{2}}[\operatorname{tr}(w^{T}P_{1}(I-\hat{P}_{12})(I-P)w^{\prime})]$
Now, using the definition of $P$ and $\hat{P}_{12}$, we have
$\mathbb{E}[w^{T}P^{T}_{01}P_{03}w^{\prime}]=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}[\operatorname{tr}(w^{T}P_{1}(I-\hat{P}_{12})w^{\prime})]$
$=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}[\operatorname{tr}(w^{T}P_{1}w^{\prime})]$
$+\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}[\operatorname{tr}(w^{T}P_{1}X_{2}(X^{T}_{2}(I-P_{1})X_{2})^{-1}X^{T}_{2}(I-P_{1})w^{\prime})]$
@@
$\mathbb{E}[w^{T}P^{T}_{01}P_{03}w^{\prime}]=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}[\operatorname{tr}(w^{T}P_{1}w^{\prime})]$
$=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\times\frac{1}{4}\mathbb{E}[(w+w^{\prime})^{T}P_{1}(w+w^{\prime})-(w-w^{\prime})^{T}P_{1}(w-w^{\prime})]$
@@
C.2 Proof of Theorems

### v2

@@ A.1 Useful Lemmas
    Without loss of generality, we can focus on finding $\mathbb{E}[\|P_{1}u\|^{2}]$ where $u\in\mathbb{R}^{p}$ and $\|u\|^{2}=1$, since $X_{1}$ is a Gaussian matrix with rank $n_{1}$, and $P_{1}=X_{1}(X_{1}^{\top}X_{1})^{-1}X^{\top}_{1}$ is an orthogonal projection matrix with a similar rank. Due to the rotational invariance of the standard normal distribution, we can assume that $P_{1}$ is a fixed matrix and instead, $u$ is a random vector uniformly sampled from the unit sphere in $\mathbb{R}^{p}$. Using the rotational invariance again, we may assume without the loss of generality that $P_{1}$ is the coordinate projection onto the first $n_{1}$ coordinates in $\mathbb{R}^{p}$. Then it holds that:
@@
    $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}[\operatorname{tr}(w^{\top}P_{1}X_{2}(X_{2}^{\top}(I-P_{1})X_{2})^{-1}X^{\top}_{2}P_{1}w)]$
    $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\operatorname{tr}(\mathbb{E}_{X_{2}}[X^{\top}_{2}P_{1}ww^{\top}P_{1}X_{2}(X_{2}^{\top}(I-P_{1})X_{2})^{-1}])]$
@@
    $\Rightarrow\mathbb{E}[\|P_{0|1}w\|^{2}]$ $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\operatorname{tr}(\mathbb{E}_{X_{2}}[X^{\top}_{2}P_{1}ww^{\top}P_{1}X_{2}]\mathbb{E}_{X_{2}}[(X_{2}^{\top}(I-P_{1})X_{2})^{-1}])]$
    $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\operatorname{tr}(\operatorname{tr}(P_{1}ww^{\top}P_{1})\mathbb{E}_{X_{2}}[(X_{2}^{\top}(I-P_{1})X_{2})^{-1}])]$
    $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\|P_{1}w\|^{2}\operatorname{tr}(\mathbb{E}_{X_{2}}[(X_{2}^{\top}(I-P_{1})X_{2})^{-1}])]$
    We focus on finding the inner expectation first. Notice that for a fixed $X_{1}$, $I-P_{1}$ is an orthogonal projection matrix with rank $p-n_{1}$. Due to the rotational invariance of the standard normal distribution, we may assume without loss of generality that $I-P_{1}$ is the projection matrix that projects onto the first $p-n_{1}$ coordinates. With this in mind, $(X_{2}^{\top}(I-P_{1})X_{2})^{-1}$ follows an inverse-Wishart distribution with an identity scale matrix $I_{n_{2}\times n_{2}}\in\mathbb{R}^{n_{2}}$ and $p-n_{1}$ degrees of freedom.
    $\Rightarrow\mathbb{E}[\|P_{0|1}w\|^{2}]$ $=\frac{n_{1}}{p}\|w\|^{2}+\mathbb{E}_{X_{1}}[\|P_{1}w\|^{2}\operatorname{tr}(\frac{I_{n_{2}\times n_{2}}}{p-n_{1}-n_{2}-1})]$
@@
Lemma A.8.
Assume matrices $X_{1}\in\mathbb{R}^{p\times n_{1}}$, $X_{2}\in\mathbb{R}^{p\times n_{2}}$ and $X_{3}\in\mathbb{R}^{p\times n_{3}}$ to be random matrices with entries being independently sampled from the standard Gaussian distribution $\mathcal{N}(0,1)$. Assume $p>n_{1}+n_{2}+n_{3}+1$ and let $X_{1:3}\in\mathbb{R}^{p\times(n_{1}+n_{2}+n_{3})}$ be the result of concatenation of $X_{1}$, $X_{2}$, $X_{3}$. Also let $X_{0|1},X_{0|2},X_{0|3}\in\mathbb{R}^{p\times(n_{1}+n_{2}+n_{3})}$ be the result of concatenating $X_{1}$, $X_{2}$ and $X_{3}$ with zero matrices such that $X_{0|1}=\begin{bmatrix}X_{1}&0&0\end{bmatrix}$ , $X_{0|2}=\begin{bmatrix}0&X_{2}&0\end{bmatrix}$ and $X_{0|3}=\begin{bmatrix}0&0&X_{3}\end{bmatrix}$. Define $P_{0|1}=X_{1:3}(X^{\top}_{1:3}X_{1:3})^{-1}X_{0|1}^{\top}$ and $P_{0|3}=X_{1:3}(X^{\top}_{1:3}X_{1:3})^{-1}X_{0|3}^{\top}$. Assuming $w,w^{\prime}\in\mathbb{R}^{p}$ are fixed given vectors, it holds that
$\mathbb{E}[w^{\top}P^{\top}_{0|1}P_{0|3}w^{\prime}]=-\frac{n_{1}n_{3}}{p(p-(n_{1}+n_{2}+n_{3})-1)}\langle w,w^{\prime}\rangle$
@@
Based on Lemma A.4, we have
$\mathbb{E}[w^{\top}P^{\top}_{0|1}P_{0|3}w^{\prime}]$ $=-\mathbb{E}[w^{\top}P_{1}(I-\hat{P}_{1:2})X_{3}(X_{3}^{\top}(I-P_{1})(I-\hat{P}_{1:2})X_{3})^{-1}X^{\top}_{3}w^{\prime}]$
$=-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[w^{\top}P_{1}(I-\hat{P}_{1:2})X_{3}(X_{3}^{\top}(I-P_{1})(I-\hat{P}_{1:2})X_{3})^{-1}X^{\top}_{3}w^{\prime}]]$
where $\hat{P}_{1:2}=X_{2}(X^{\top}_{2}(I-P_{1})X_{2})^{-1}X^{\top}_{2}(I-P_{1})$. Now denote $P=(I-P_{1})(I-\hat{P}_{1:2})$. With a fixed $X_{1}$ and $X_{2}$, $P$ is an orthogonal projection matrix, meaning that $P^{\top}=P^{2}=P$. Therefore,
$\mathbb{E}[w^{\top}P^{\top}_{0|1}P_{0|3}w^{\prime}]$ $=-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[w^{\top}P_{1}(I-\hat{P}_{1:2})X_{3}(X_{3}^{\top}PX_{3})^{-1}X^{\top}_{3}(P+I-P)w^{\prime}]]$
$=-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[w^{\top}P_{1}(I-\hat{P}_{1:2})X_{3}(X_{3}^{\top}PX_{3})^{-1}X^{\top}_{3}Pw^{\prime}]]$
$-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[w^{\top}P_{1}(I-\hat{P}_{1:2})X_{3}(X_{3}^{\top}PX_{3})^{-1}X^{\top}_{3}(I-P)w^{\prime}]]$
@@
$P_{1}(I-\hat{P}_{1:2})X_{3}$ $\operatorname{\perp\!\!\!\perp}(I-P_{1})(I-\hat{P}_{1:2})X_{3}$
@@
$\mathbb{E}[w^{\top}P^{\top}_{0|1}P_{0|3}w^{\prime}]$ $=-\mathbb{E}_{X_{1},X_{2}}[w^{\top}P_{1}\mathbb{E}_{X_{3}}[(I-\hat{P}_{1:2})X_{3}]\mathbb{E}_{X_{3}}[(X_{3}^{\top}PX_{3})^{-1}X^{\top}_{3}Pw^{\prime}]]$
$-\mathbb{E}_{X_{1},X_{2}}[\mathbb{E}_{X_{3}}[\operatorname{tr}(X^{\top}_{3}(I-P)w^{\prime}w^{\top}P_{1}(I-\hat{P}_{1:2})X_{3}(X_{3}^{\top}PX_{3})^{-1}])]$
$=0-\mathbb{E}_{X_{1},X_{2}}[\operatorname{tr}(\mathbb{E}_{X_{3}}[X^{\top}_{3}(I-P)w^{\prime}w^{\top}P_{1}(I-\hat{P}_{1:2})X_{3}]\mathbb{E}_{X_{3}}[(X_{3}^{\top}PX_{3})^{-1}])]$
$=-\mathbb{E}_{X_{1},X_{2}}[\operatorname{tr}((I-P)w^{\prime}w^{\top}P_{1}(I-\hat{P}_{1:2}))\operatorname{tr}(\mathbb{E}_{X_{3}}[(X_{3}^{\top}PX_{3})^{-1}])]$
Here we again use a technique similar to the one we used in Lemma A.7 part (iii) by first focusing on the inner expectation. Assume a fixed $X_{1}$ and $X_{2}$. Then $I-P_{1}$ and $P_{1}$ are orthogonal projections with ranks $p-n_{1}$ and $n_{1}$ respectively. Additionally, $(I-P_{1})\hat{P}_{1:2}$ is an orthogonal projection matrix with rank $n_{2}$ and $p-n_{1}>n_{2}$. On the other hand, $P_{1}$ and $(I-P_{1})\hat{P}_{1:2}$ are orthogonal projections. Therefore, $P_{1}+(I-P_{1})\hat{P}_{1:2}$ is a projection matrix with rank $n_{1}+n_{2}$ and $P=I-(P_{1}+(I-P_{1})\hat{P}_{1:2})$ is a projection matrix with rank $p-(n_{1}+n_{2})$. When taking the inner expectation only w.r.t. $X_{3}$, due to the rotational invariance property, we may assume that $(I-P_{1})(I-\hat{P}_{1:2})$ is the coordinate projection onto the first $p-(n_{1}+n_{2})$ coordinates. Therefore, $(X_{3}^{\top}(I-P_{1})(I-\hat{P}_{1:2})X_{3})^{-1}\sim\mathcal{IW}(I_{n_{3}\times n_{3}},p-(n_{1}+n_{2}))$ and we have:
$\Rightarrow\mathbb{E}[w^{\top}P^{\top}_{0|1}P_{0|3}w^{\prime}]$ $=-\mathbb{E}_{X_{1},X_{2}}[\operatorname{tr}((I-P)w^{\prime}w^{\top}P_{1}(I-\hat{P}_{1:2}))\operatorname{tr}(\frac{I_{n_{3}\times n_{3}}}{p-(n_{1}+n_{2})-n_{3}-1})]$
$=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}_{X_{1},X_{2}}[\operatorname{tr}(w^{\top}P_{1}(I-\hat{P}_{1:2})(I-P)w^{\prime})]$
Now, using the definition of $P$ and $\hat{P}_{1:2}$, we have
$\mathbb{E}[w^{\top}P^{\top}_{0|1}P_{0|3}w^{\prime}]=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}[\operatorname{tr}(w^{\top}P_{1}(I-\hat{P}_{1:2})w^{\prime})]$
$=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}[\operatorname{tr}(w^{\top}P_{1}w^{\prime})]$
$+\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}[\operatorname{tr}(w^{\top}P_{1}X_{2}(X^{\top}_{2}(I-P_{1})X_{2})^{-1}X^{\top}_{2}(I-P_{1})w^{\prime})]$
@@
$\mathbb{E}[w^{\top}P^{\top}_{0|1}P_{0|3}w^{\prime}]=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\mathbb{E}[\operatorname{tr}(w^{\top}P_{1}w^{\prime})]$
$=-\frac{n_{3}}{p-(n_{1}+n_{2}+n_{3})-1}\times\frac{1}{4}\mathbb{E}[(w+w^{\prime})^{\top}P_{1}(w+w^{\prime})-(w-w^{\prime})^{\top}P_{1}(w-w^{\prime})]$
@@
As a concluding remark to this section, notice that When $X$ is a Gaussian matrix with iid entries, $X^{T}X$ is almost surely invertible (i.e. the probability measure of $X^{T}X$ not being invertible is zero). Therefore, in what comes next, we assume that all Gaussian matrices in this form are invertible without loss of generality.
A.2 Proof of Theorems